A winery has a vat with two pipes leading to it. The inlet pipe can fill the vat in 3 hours, while the outlet pipe can empty it in 7 hours. How long will it take to fill the Vat with both pipes are left open?
Answer by richwmiller(17219)

1/3-1/7=1/x
7/21-3/21=1/x
4/21=1/x
x=21/4=5.25 hours